A volume of 225 mL of hydrogen is collected over water; the water level in the collecting vessel is the same as the outside level. Atmospheric pressure is 756.0 Torr and the temperature is 25 °C. The vapor pressure of water at various temperatures can be found in this table. Calculate the molar mass of the metal.
A 0.495 gram sample of a metal, M, reacts completely with sulfuric acid according to:
1
Expert's answer
2018-12-11T02:18:10-0500

M(s) + H2SO4(aq) → MSO4(aq) + H2(g)


First, you need to get to the moles of H2 produced. Subtract the vapor pressure of H2O at 25 °C from the pressure of the gas: 760.0 - 23.8 = 736.2torr/760 = 0.969 atm


Now, use the ideal gas law to calculate moles of H2:


PV = nRT


0.969 atm*(0.225 L) = n*(0.0821 Latm/molK)*(298 K)


n = 8.91*10^-3 mol H2


Since the stoichiometry of the reaction is 1:1, you had in the sample 8.91*10^-3 mol of metal M(s). So, the molar mass of M is:


0.495 g / 8.91*10^-3 mol = 55.89 g/mol


The metal was most likely Fe.